Required :
(a) What is the profitability index for the cashflows if the relevant discount rate is 10 percent?

  

(b) What is the profitability index for the cashflows if the relevant discount rate is 14 percent?

  

(c) What is the profitability index for the cashflows if the relevant discount rate is 22 percent?
Year Cash Flow
0 −$8,400          
1 3,000          
2 5,000          
3 4,000          
Answer #1

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

At 10%:

Present value of inflows=3000/1.1+5000/1.1^2+4000/1.1^3

=$9864.763336

PI=Present value of inflows/Present value of outflows

=$9864.763336/$8400

=1.17(Approx)

At 14%:

Present value of inflows=3000/1.14+5000/1.14^2+4000/1.14^3

=$9178.802655

PI=Present value of inflows/Present value of outflows

=$9178.802655/$8400

=1.09(Approx)

At 22%:

Present value of inflows=3000/1.22+5000/1.22^2+4000/1.22^3

=$8021.155956

PI=Present value of inflows/Present value of outflows

=$8021.155956/$8400

=0.95(Approx).